OpenStudy (anonymous):

Solve: 7^(x – 5) = 74^(x + 7)

OpenStudy (anonymous):

logariddims.

OpenStudy (anonymous):

start with \[(x-5)\ln(7)=(x+7)\ln(74)\] and then to a bunch of algebra to solve for x

OpenStudy (anonymous):

distribute, collect like terms etc

OpenStudy (anonymous):

what did you get for an answer?

OpenStudy (anonymous):

xln7 - 5ln7 = xln74 + 7ln74 xln7 - xln74 = 4ln7 + 7ln74 x(ln7 - ln74) = 4ln7 + 7ln74 x = (4ln7 + 7ln74) / (ln7 - ln74)